OpenStudy (anonymous):

one What is the equation in point-slope form of the line passing through (0, 2) and (1, 5)?

OpenStudy (texaschic101):

first we need to find the slope by using the slope formula : slope(m) = (y2 - y1) / (x2 - x1) (0,2) x1 = 0 and y1 = 2 (1,5) x2 = 1 and y2 = 5 now lets sub our info into the formula... m = (5 - 2) / (1 - 0) m = 3/1 or just 3 do you understand so far ?

OpenStudy (anonymous):

Y es I got y-5= 3(x-)

OpenStudy (texaschic101):

point slope formula is y - y1 = m(x - x1) slope = 3 as for the points, you can use either set of points... using points (0,2) you get : y - 2 = 3(x - 0) using points (1,5) you get : y - 5 = 3(x - 1) understand ?

OpenStudy (anonymous):

Yes i do thankyou!

OpenStudy (texaschic101):

good to hear :)